# Criticality at work: scaling in the mouse cortex enhances performance

### v1

@@ 1.1 Data
The data used in this study comes from the Allen Visual Behavior Neuropixels dataset [50], which includes neuronal activity recorded from the V1 region of mice using Neuropixels 1.0 probes. Each recording spans approximately 2 hours and 25 minutes. Prior to the experiment, mice were trained to lick a device upon detecting an image change in order to receive a reward. During the first hour of the experiment, they were shown a series of images under these trained conditions. Following this phase, mice were shown Gabor patches for 15 minutes, then exposed to black and white flashes, and finally shown a grey screen for 5 minutes. In the third phase, the natural images are shown a second time in the same order as before, but this time the lick apparatus is retracted. The raw electrophysiological data was pre-processed and sorted into spike trains using the Kilosort2 algorithm [47]. We selected all 62 sessions from the dataset that had more than 128 recorded units, which included both single-unit activity (SUA) and multi-unit activity (MUA). The data was then divided into these three experimental phases, binned into 50-ms intervals, and segmented into 30-s windows for analysis.
@@
$CV=\frac{\sigma}{\mu},$ (1)
where $\mu$ and $\sigma$ are the mean and the standard deviation of the activity across the window. Windows where $CV>5$ were excluded, as such high values are not typically observed in awake mice and likely result from the chosen window size, indicating artefacts.
1.2 Surrogate data
Throughout our analysis, we compared our results with surrogate data. This surrogate data was generated by shuffling the inter-spike intervals of each unit within each time window, effectively breaking correlations between units and producing trivial kurtosis values and exponents for comparison.
1.3 Phenomenological Renormalisation Group (PRG) analysis
1.3.1 Momentum-space renormalisation and kurtosis
@@
where $\phi_{i}$ is the time series activity of unit $i$. Taking its ranked eigenvalues ${\lambda_{1}>\lambda_{2}>\ldots>\lambda_{N}}$ and the components $u_{\mu i}$ of their respective eigenvectors, such that
@@
where $\mathcal{Z}_{i}(N_{\mathrm{cutoff}})$ is such that $\text{var}(\psi_{i})=1$, as shown in Fig. 1.d for $N_{\mathrm{cutoff}}=N$, $N/2$, $N/4$, $N/8$, $N/16$ [48, 45].
From the distribution of $\phi_{i}$ at the final coarse graining step ($N_{\mathrm{cutoff}}=N/16$) we calculate the kurtosis $\kappa=\langle\psi^{4}\rangle/\langle\psi^{2}\rangle^{2}$ (Fig. 1.f). Non critical dynamics converge towards a Gaussian distribution ($\kappa=3$) as we reduce the number of eigenvectors involved in the construction of the projector (4), while critical dynamics does not. Note that coarse-grained distributions do not converge perfectly to a Gaussian even for surrogate data (with lowest kurtoses around $\kappa\simeq 6$), due to the finite number of bins within a window.
1.3.2 Real-space renormalisation and exponents
In order to obtain the PRG scaling exponents presented in Fig. 3.(b-e) we use a different phenomenological renormalisation process, to which we refer as real-space renormalisation [37]. In this analysis we cluster together the activities of the most correlated pair of neurons repeatedly until we exhaust all neurons:
@@ 1.3.2 Real-space renormalisation and exponents
where $j$ is the most correlated neuron to $i$. This procedure is repeated $k$ times, after which we have $N/2^{k}$ coarse-grained units (or “clusters”), each comprising the sum of $K=2^{k}$ original spiking units. We follow some quantities that are expected to scale with non-trivial power laws only close to the critical regime.
@@
The spectrum of eigenvalues of the covariance matrix (Eq. 2) when ordered in a descending order should follow
@@
For trivial data, this scaling is broken and $\mu$ is undefined.
@@
1.4 Hit rate, trial shuffling and relative performance
1.4.1 Hit rate
During the first phase of the Allen Brain Observatory Visual Behavior Neuropixels experiment, mice performed a visual recognition task. To assess the impact of criticality on performance, we subdivided the data into 30-s intervals. After applying the PRG procedure streamlined in Methods 1.3, intervals with a kurtosis above a chosen $\kappa_{th}$ were labelled as scaling windows. We then calculated the hit rate for these scaling windows by summing all correctly performed tasks (hits) within them and dividing by the total number of tasks in these intervals.
$\mathrm{Hit~{}rate(Scaling)}$ $=\frac{\mathrm{Total~{}number~{}of~{}sucessful~{}tasks(windows~{}with~{}\kappa>\kappa{th})}}{\mathrm{Total~{}number~{}of~{}tasks(windows~{}with~{}\kappa>\kappa{th})}},$ (11)
$\mathrm{Hit~{}rate(No~{}scaling)}$ $=\frac{\mathrm{Total~{}number~{}of~{}sucessful~{}tasks(windows~{}with~{}\kappa<\kappa{th})}}{\mathrm{Total~{}number~{}of~{}tasks(windows~{}with~{}\kappa<\kappa{th})}}.$ (12)
The standard error of the mean was estimated through a bootstrapping statistic method, by choosing $\mathcal{N}$ samples with replacement from the subset of time windows above (below) the kurtosis threshold, averaging and taking the standard deviation over $n=10^{5}$ repetitions.
@@ 1.4.1 Hit rate
$\mathrm{Relative~{}performance}=\frac{\mathrm{Hit~{}rate(Scaling)}-\mathrm{Hit~{}rate(No~{}scaling)}}{\mathrm{Hit~{}rate(No~{}scaling)}}.$ (13)
1.4.2 Trial shuffled data
@@
To control for artifacts, we generated a trial-shuffled version of the data. For each trial, we randomly selected another trial within the same session and swapped their outcomes (hit or miss, Fig. 2.b). This process preserves the total number of hits and misses within each session, as well as the total number of trials within each window.
@@
For the trial-shuffled data, we calculated the hit rate for the scaling windows as described in Methods 1.4.1, repeating this procedure 500 times. The mean and standard deviation across these repetitions are presented in (Fig. 2.c).
1.5 Statistical procedures
All the results are expressed as mean $\pm$ standard deviation. For statistical comparisons between distributions we used Mann-Whitney U test. We use the convention for representing $p$-values: $p\geq 0.05$ (n.s.), $p<0.05$ (*), $p<0.01$ (**), $p<0.001$ (***), where “n.s.” stands for not significant. The p-value for the comparisons stated in the text are the following:
  • $\alpha_{1}$ vs $\alpha_{2}$p-value$=8\times 10^{-5}$
  • $\alpha_{2}$ vs $\alpha_{3}$p-value$=4\times 10^{-19}$,
  • $\alpha_{1}$ vs $\alpha_{3}$p-value$=8\times 10^{-21}$,
  • $\beta_{1}$ vs $\beta_{2}$p-value$=0.17$,
  • $\beta_{2}$ vs $\beta_{3}$p-value$=3\times 10^{-10}$,
  • $\beta_{1}$ vs $\beta_{3}$p-value$=3\times 10^{-13}$,
  • $z_{1}$ vs $z_{2}$p-value$=6\times 10^{-5}$,
  • $z_{2}$ vs $z_{3}$p-value$=6\times 10^{-6}$,
  • $z_{1}$ vs $z_{3}$p-value$=3\times 10^{-10}$,
  • $\mu_{1}$ vs $\mu_{2}$p-value$=4\times 10^{-13}$,
  • $\mu_{2}$ vs $\mu_{3}$p-value$=3\times 10^{-8}$,
  • $\mu_{1}$ vs $\mu_{2}$p-value$=3\times 10^{-8}$.
[…]
Supplementary information
This article contains supplementary material.
[…]
Acknowledgements The authors acknowledge funding from the Brazilian agencies CNPq (Grants No. 140660/2022-4, 308703/2022-7 and 314094/2023-7), FACEPE (Grant No. APQ-1187-1.05/22) and FINEP (Grant No. NeuroAssist 01.24.0124.00).
Declarations
Funding
The authors acknowledge funding from the Brazilian agencies CNPq (Grants No. 140660/2022-4, 308703/2022-7 and 314094/2023-7) and FACEPE (Grant No. APQ-1187-1.05/22) and FINEP (Grant No. NeuroAssist 01.24.0124.00).
Competing interests
Authors declare no competing interests.
Ethics approval and consent to participate
Not applicable.
Consent for publication
All authors consent to the publication.
Data availability
Data from all the 62 sessions used are available to download via the AllenSDK development kit at https://allensdk.readthedocs.io/.
Materials availability
Not applicable.
Code availability
The code used can be made available upon request.
Author contribution
MC and DMC designed the study. DMC provided initial code for the analysis. GGC refined the code, ran the analysis and made the plots. All authors discussed the results and wrote the paper.
Appendix A Supplementary Information
A.1 Effects of different time window lengths
Our choice of time window length to perform the analysis was made to balance two competing considerations: on the one hand, to have a sufficiently long time for the surrogate procedure to effectively shuffle the data within each window; on the other hand, to avoid an overly broad time frame that could blend different cortical dynamics.
To examine the sensitivity of our results to this choice of window length, we repeated the analyses using 10-s and 60-s windows. This allowed us to evaluate the impact of shorter and longer time windows on our performance results.
Figure 4: Relation of task performance and kurtosis for different windows sizes. (a-b), hit rates for the 10-s and 60-s windows exhibit a similar trend to those of 30-s windows, with increasing hit rates as the threshold ($\kappa_{th}$) is increased. (c-d), Relative performance follows a similar pattern, deviating from zero around $\kappa\approx 7$.
We observe that the change in window size choice does not qualitatively alter our results (Fig. 4). We consistently observe that windows with higher kurtosis are associated with improved performance.
A.2 Presence of noise units
A natural question arises regarding whether our choice to use a greedy spike sorting approach (potentially including noise units) could influence the results. To address this, we repeated the analyses using only the 43 sessions that retained more than 128 units after applying more strict sorting criteria, while maintaining the 30-s time window.
Figure 5: Relation of task performance and kurtosis for only “no noise” units. a-b We repeat the plots presented in the main text, the results of increased performance are kept.
The results remained qualitatively consistent with those observed in the initial analyses (Fig. 5). This outcome is expected, as a procedure in the spirit of the renormalisation group, which captures large-scale properties, should be resilient to the inclusion of noise units.
A.3 Exponents vs. threshold
The exponents presented in (Fig. 3) are calculated across all windows of a given session. However, it is reasonable to question this approach, as only a subset of these windows may truly exhibit scaling behaviour. To address this, we analyse how the exponents $\alpha$, $\beta$, $\mu$ and $z$ change when calculated using only the windows with $\kappa>\kappa_{th}$.
Figure 6: Exponents variation with $\kappa_{th}$. (a-d) Exponent values show only minimal variation with different $\kappa_{th}$. The light solid lines represent the surrogate data, while dashed lines show the kurtosis distribution $P(\kappa)$ of each experimental phase.
We find that the exponent values vary only slightly and remain within their respective error bars (Fig. 6). The larger fluctuations for larger kurtosis values are due to the small number of windows (note that $P(\kappa)$ has decreased considerably in those regions).
A.4 Activity kurtosis vs. coarse-grained activity kurtosis
One could wonder whether coarse-graining the neural activity is necessary to compute the kurtosis, or if examining the kurtosis of the original, non-coarse-grained activity distribution would suffice. Our analysis shows that the initial kurtosis $\kappa_{0}$ is a very poor predictor of the coarse-grained kurtosis $\kappa_{final}$ (Fig. 7).
Figure 7: Comparison of initial and final kurtoses. For each window we plot the value of the kurtosis of the activity distribution after the coarse-graining ($\kappa_{\mathrm{final}}$) against their initial kurtosis from the activity distribution ($\kappa_{0}$). There is no clear tendency in the relation between those two variables.

### v2

@@ Acknowledgements.
The data used in this study come from the Allen Visual Behavior Neuropixels dataset [60], which includes neuronal activity recorded from the V1 region of mice using Neuropixels 1.0 probes. Each recording spans approximately 2 h and 25 min. Prior to the experiment, mice were trained to lick a device upon detecting an image change in order to receive a reward. During the first hour of the experiment, they were shown a series of images under these trained conditions. Following this phase, mice were shown Gabor patches for 15 min, then exposed to black and white flashes, and finally shown a gray screen for 5 min. In the third phase, the natural images are shown a second time in the same order as before, but this time the lick apparatus is retracted. The raw electrophysiological data were preprocessed and sorted into spike trains using the Kilosort2 algorithm [52]. We selected all 62 sessions from the dataset that had more than 128 recorded units, which included both single-unit activity (SUA) and multiunit activity (MUA). The data were then divided into these three experimental phases, binned into 50-ms intervals, and segmented into 30-s windows for analysis.
@@
$\textrm{CV}=\frac{\sigma}{\mu},$ (1)
where $\mu$ and $\sigma$ are the mean and the standard deviation of the activity across the window. Windows where $\textrm{CV}>5$ were excluded, as such high values are not typically observed in awake mice and likely result from the chosen window size, indicating artefacts.
.2 Surrogate data
Throughout our analysis, we compared our results with surrogate data. These surrogate data were generated by shuffling the interspike intervals of each unit within each time window, effectively breaking correlations between units and producing trivial kurtosis values and exponents for comparison.
.3 Momentum-space renormalization and kurtosis
@@
where $\phi_{i}$ is the time series activity of unit $i$. Taking its ranked eigenvalues $\lambda_{1}>\lambda_{2}>\cdots>\lambda_{N}$} and the components $u_{\mu i}$ of their respective eigenvectors, such that
@@
where $\mathcal{Z}_{i}(N_{\mathrm{cutoff}})$ is such that $\text{var}(\psi_{i})=1$, as shown in Fig. 1D for $N_{\mathrm{cutoff}}=N$, $N/2$, $N/4$, $N/8$, $N/16$ [54, 50].
From the distribution of $\phi_{i}$ at the final coarse graining step ($N_{\mathrm{cutoff}}=N/16$), we calculate the kurtosis $\kappa=\langle\psi^{4}\rangle/\langle\psi^{2}\rangle^{2}$ [Fig. 1F]. Non-critical dynamics converges towards a Gaussian distribution ($\kappa=3$) as we reduce the number of eigenvectors involved in the construction of the projector [Eq. 4], while critical dynamics does not. Note that coarse-grained distributions do not converge perfectly to a Gaussian even for surrogate data (with the lowest kurtoses around $\kappa\simeq 6$), due to the finite number of bins within a window.
.4 Real-space renormalization and exponents
To obtain the PRG scaling exponents presented in Figs. 4B-4E we use a different phenomenological renormalization process, which we call real-space renormalization [42]. In this analysis we cluster together the activities of the most correlated pair of neurons repeatedly until we exhaust all neurons:
@@
where $j$ is the most correlated neuron to $i$. This procedure is repeated $k$ times, after which we have $N/2^{k}$ coarse-grained units (or “clusters”), each comprising the sum of $K=2^{k}$ original spiking units. We follow some quantities that are expected to scale with nontrivial power laws only close to the critical regime.
@@
The spectrum of eigenvalues of the covariance matrix [Eq. 2] when ordered in descending order should follow
@@
For trivial data, this scaling is broken and $\mu$ is undefined. Time windows with a fit quality of $r^{2}<0.85$ for this power-law have their corresponding exponent discarded, as it cannot be reliably considered indicative of scaling.
@@
.5 Hit rate
During the first phase of the Allen Brain Observatory Visual Behavior Neuropixels experiment, mice performed a visual recognition task. To assess the impact of criticality on performance, we subdivided the data into 30-s intervals. After applying the PRG procedure streamlined in the previous section, intervals with a kurtosis above a chosen $\kappa_{\textrm{th}}$ were labeled as scaling windows. Within each 30-s window only three to five trials are performed, leading to poor statistical estimation of performance on a window-to-window basis. For this reason we chose to calculate the hit rate for these scaling windows by summing all correctly performed tasks (hits) within them and dividing by the total number of tasks in these intervals,
$\mathrm{Hit\leavevmode\nobreak\ rate\leavevmode\nobreak\ (Scaling)}$ $=\frac{\mathrm{Total\leavevmode\nobreak\ number\leavevmode\nobreak\ of\leavevmode\nobreak\ sucessful\leavevmode\nobreak\ tasks\leavevmode\nobreak\ (windows\leavevmode\nobreak\ with\leavevmode\nobreak\ \kappa>}\kappa_{\textrm{th}})}{\mathrm{Total\leavevmode\nobreak\ number\leavevmode\nobreak\ of\leavevmode\nobreak\ tasks\leavevmode\nobreak\ (windows\leavevmode\nobreak\ with\leavevmode\nobreak\ \kappa>}\kappa_{\textrm{th}})},$ (11)
$\mathrm{Hit\leavevmode\nobreak\ rate\leavevmode\nobreak\ (No\leavevmode\nobreak\ scaling)}$ $=\frac{\mathrm{Total\leavevmode\nobreak\ number\leavevmode\nobreak\ of\leavevmode\nobreak\ sucessful\leavevmode\nobreak\ tasks\leavevmode\nobreak\ (windows\leavevmode\nobreak\ with\leavevmode\nobreak\ \kappa<}\kappa{{}_{\textrm{th}}})}{\mathrm{Total\leavevmode\nobreak\ number\leavevmode\nobreak\ of\leavevmode\nobreak\ tasks\leavevmode\nobreak\ (windows\leavevmode\nobreak\ with\leavevmode\nobreak\ \kappa<}\kappa_{\textrm{th}})}.$ (12)
The standard error of the mean was estimated through a bootstrapping statistic method by choosing $\mathcal{N}$ samples with replacement from the subset of time windows above (below) the kurtosis threshold, averaging and taking the standard deviation over $n=10^{5}$ repetitions.
@@
$\mathrm{Relative\leavevmode\nobreak\ performance}=\frac{\mathrm{Hit\leavevmode\nobreak\ rate\leavevmode\nobreak\ (Scaling)}-\mathrm{Hit\leavevmode\nobreak\ rate\leavevmode\nobreak\ (No\leavevmode\nobreak\ scaling)}}{\mathrm{Hit\leavevmode\nobreak\ rate\leavevmode\nobreak\ (No\leavevmode\nobreak\ scaling)}}.$ (13)
.6 Trial-shuffled data
@@
To control for artifacts, we generated a trial-shuffled version of the data. For each trial, we randomly selected another trial within the same session and swapped their outcomes [hit or miss, Fig. 3B]. This process preserves the total number of hits and misses within each session, as well as the total number of trials within each window.
@@
For the trial-shuffled data, we calculated the hit rate for the scaling windows as described in the previous section, repeating this procedure 500 times. The mean and standard deviation across these repetitions are presented in Fig. 3C.